OpenStudy (anonymous):

The area of the trapezoid is 24 and the bases are 3 and 5. Find the height.

OpenStudy (anonymous):

\[A = \frac h2(b_{1} + b_{2})\] \[24 = \frac h2(3 + 5)\] \[24 = \frac h2(8)\] \[3 = \frac h2\] \[h = 6\]